OpenStudy (anonymous):

Help! What's the Differential dy of the function y = 1/3cos(6πx-1/2)

OpenStudy (anonymous):

(1/3) (-sin(6 pi x))6pi=dy/dx dy=(1/3) (-sin(6 pi x))6pi dx

OpenStudy (anonymous):

thanks